OpenStudy (anonymous):

@kayders1997 72 = 32q^2

OpenStudy (anonymous):

I have to solve this but it needs to have two answers

OpenStudy (kayders1997):

Do you have an idea of where to start?

OpenStudy (anonymous):

set to equal zero!

OpenStudy (kayders1997):

You don't have to in this one

OpenStudy (anonymous):

\[0 = 32q ^{2} -72\]

OpenStudy (kayders1997):

You could isolate to get q^2 on one side

OpenStudy (kayders1997):

You could do that too and factor out the greatest common factor I suppose :)

OpenStudy (kayders1997):

Which way do you want to do it :)

OpenStudy (anonymous):

the way that I set it up if we could please

OpenStudy (kayders1997):

Yeah we can do it that way! I can show you my way too after we're done if you would like

OpenStudy (kayders1997):

So what is the greatest common factor between the 2

OpenStudy (anonymous):

8?

OpenStudy (kayders1997):

Yes I think that is it, so how would that look after you take that out?

OpenStudy (anonymous):

\[0 = 8 (4q ^{2} - 9)\] ?

OpenStudy (kayders1997):

Good now you have to split up the two parts do you know how to do that?

OpenStudy (kayders1997):

Where the first part will equal 0 and the second part will equal 0

OpenStudy (anonymous):

\[0 = 8(2q -3) (2q-3)\] ?

OpenStudy (anonymous):

I am not entirely sure of what i'm doing

OpenStudy (kayders1997):

You don't have to factor again

OpenStudy (kayders1997):

Since we have the outside of 8 we can set that equal to 0 so 8=0 is that true though?

OpenStudy (anonymous):

no

OpenStudy (kayders1997):

good, now we can set the other part equal to 0 so 4q^2-9=0 the reason why we can do this is because we have the outside part the part where you factored something out and an inside part

OpenStudy (kayders1997):

So 4q^2-9=0 do you know how to do this part?

OpenStudy (anonymous):

sooo just ignore the 8?

OpenStudy (kayders1997):

Yes because 8 does not equal 0

OpenStudy (anonymous):

alright. moving on then. \[4q ^{2}=9\] Is that the next step?

OpenStudy (kayders1997):

Yes good

OpenStudy (kayders1997):

Now do you know what's next? :)

OpenStudy (anonymous):

something having to do with perfect squares right?

OpenStudy (kayders1997):

You don't have to do that think of PEMDAS backwards we just want q^2=some number

OpenStudy (anonymous):

so division

OpenStudy (kayders1997):

good yes :) so you get?

OpenStudy (anonymous):

\[q^{2} = \frac{ 9 }{ 4 }\] Is that the answer?

OpenStudy (kayders1997):

Almost!!!

OpenStudy (anonymous):

it doesn't seem quite right...

OpenStudy (kayders1997):

But that is the correct step

OpenStudy (kayders1997):

Now how do you undo something that is being squared?

OpenStudy (anonymous):

I know what it is but I can't for the life of me remember what it's called.

OpenStudy (anonymous):

OH WAIT

OpenStudy (anonymous):

THE ROOT

OpenStudy (kayders1997):

:)

OpenStudy (anonymous):

IS IT THE ROOT

OpenStudy (kayders1997):

Yes square root

OpenStudy (kayders1997):

So...you get!!!!! :)

OpenStudy (anonymous):

q= \[q = \frac{ 3 }{ 2 }\]

OpenStudy (kayders1997):

Yes :) and remember when you take the square root you get a plus minus answer so it is q=+3/2 and q=-3/2 and we could check that by putting those back into the equation

OpenStudy (kayders1997):

Do you wanna see the way I approached it?

OpenStudy (anonymous):

wait a second. break that last bit down again please? it can be a positive or negative answer right? so in order to find out you have to send the fraction through what equation to see if the fraction (not the individual numbers) is + or - ?

OpenStudy (anonymous):

brb getting food

OpenStudy (anonymous):

back

OpenStudy (kayders1997):

Okay so when you take the square root you will always get a + answer and a - answer, it's just a rule like if you think about the graph x^2 where is two x values when the y value is 1 x=1 and x=-1 that's why, because this is an even function so it will have the same look on both sides of the graph, and than the second part we can plug in the x values back into the function 72=32q^2 so we got -3/2 so let's plug that in 72=32(-3/2)^2 when you do the (-3/2)^2 you get 2.25 and than times that by 32 you get 72 and 72=72 if you do that with the positive answer so put in 3/2 so 72=32(3/2)^2 so again you do the (3/2)^2 and you got 2.25 multiply that by 32 72=72 so that's how you can check the answers

OpenStudy (anonymous):

if it's the same both ways do you write the answer like this: \[q = \pm \frac{ 3 }{ 2 }\]

OpenStudy (kayders1997):

Yes you can

OpenStudy (kayders1997):

Or you can write it separate doesn't matter they are both correct :)

OpenStudy (kayders1997):

Do you want to see my way, it is quicker if you don't that's okay too :)

OpenStudy (anonymous):

yeah! (i may or may not understand... just a warning)

OpenStudy (kayders1997):

That's okay I'all try to explain step by step as best as I can

OpenStudy (kayders1997):

72=32q^2 first I know we wanted to solve for q so first I divided by 32 so you would only have q^2 on the one side so you would get q^2=9/4 does that make sense?

OpenStudy (anonymous):

suprisingly...yes

OpenStudy (kayders1997):

And than we just took the root like we did before and we get q=+3/2 and q=-3/2

OpenStudy (kayders1997):

Does that make sense? :)

OpenStudy (anonymous):

yeah. I can't believe my teacher doesn't use this method in class

OpenStudy (kayders1997):

Really? Omg!!! It's super quick, and in my mind pretty easy, factoring works too as we just seen, but yes there is more than one way to solve problems like these :)

OpenStudy (anonymous):

weird. did i mention that we use a common core A1 book? that might have something to do with her not using this.

OpenStudy (kayders1997):

Yes you have, I don't think it should matter though I want to become a math teacher myself, but I haven't gotten to things like common core standards for classes, I'm not even in college yet :D

OpenStudy (kayders1997):

But I wouldn't think it would matter this is a way to solve an equation for a variable which has to be a common core standard in an algebra book

OpenStudy (anonymous):